OpenStudy (anonymous):

HELP!! Determine and state if the sequence below is Arithmetic or Geometric. Then find the value of the term. 4,-8, 16, -32... Find a10

OpenStudy (anonymous):

-8-4=-12,16-(-8)=24 So it is not arithmetic \[\frac{ -8 }{ 4 }=-2,\frac{ 16 }{ -8 }=-2\] so c.r. is same calculate a10

OpenStudy (anonymous):

okay so what is A10?

OpenStudy (anonymous):

\[a_{n}=ar ^{n-1}\] a=4 r=-2 n=10 find \[a _{10}\]